OpenStudy (anonymous):

19 = h/3 - 8 what does h equal explain the answer

OpenStudy (anonymous):

What you do is add 8 to 19 to cancel it out. Now you know that h/3 = 27 You multiply 27 by 3 to get h. The answer would be 81. If that doesn't help I could put a longer explanation.

OpenStudy (sarahjoy26):

Yeah @Bamawam I totally agree.

OpenStudy (johnweldon1993):

Basically what this says in words is When 8 is subtracted from some number that is divided by 3...you get 19 So \[\large \frac{h}{3} - 8 = 19\] If we cancel out that subtraction of 8...by adding 8 to both sides...we would have \[\large \frac{h}{3} = 27\] Now...if some number divided by 3 = 27....then 27 times 3 would be that number right? \[\large h = 27 \times 3 = ?\]

OpenStudy (anonymous):

yup thanks everyone